OpenStudy (anonymous):

A point P whose x-coordinates is a is taken on the line y=3x-7. If Q is the point(4,1) show that PQ2 = 10a2-56a+80. Find the value of a which will make the expression a minimum. Hence show that the coordinates of N, the foot of the perpendicular from Q to the line are (24/5 , 1 2/5). Find the equation of QN. I did the first part I think a is 2.8 but it's the last part I can't do. Please help!

Directrix (directrix):

Is this --> PQ2 supposed to be (PQ) ^2 ? @lionely And, 10a2 --> supposed to be 10 a ^ 2 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

sorry I didn't notice

Directrix (directrix):

Do you agree with the following: Point P has coordinates (a, 3a - 7) Point Q has coordinates (4,1) (PQ)^2 = (4-a)^2 + [ 1 - (3a-7) ] ^ 2

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

PQ^2 = 10a^2 - 56a +80 dy/dx = 20a -56 a= 2.8

OpenStudy (anonymous):

sorry for awfully late reply I'm still trying to attempt other questions.

OpenStudy (anonymous):

I'm still here.

Directrix (directrix):

@lionely I think you may have to use implicit differentiation for the equation. (PQ)^2 = 10a^2 - 56a +80 can be thought of as y^2 = 10x^2 - 56x + 80 where y = PQ and a = x. y^2 = 10x^2 - 56x + 80 y = sqrt(10x^2 - 56x + 80) or y = - sqrt(10x^2 - 56x + 80) So, y' does not appear to equal dy/dx = 20a -56 http://www.wolframalpha.com/input/?i=derivative+of+y+%3D+sqrt%2810x%5E2+-+56x+%2B+80%29 I continue to think that the value of a is incorrect and that is the impasse to solving the second portion of the problem.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

So I have to square root it and then differentiate?

OpenStudy (anonymous):

In my book it's written PQ^2 the whole thing isn't squared.

hartnn (hartnn):

you had to minimize the expression 10a2-56a+80 you have correctly minimized it to get a= 2.8 now, what you minimized was the distance from a point to the line, which is also 'Perpendicular Distance' geometrically. so, the co-ordinates you get by minimizing the distance will be co-ordinates of N only. so, x-co-ordinate of N = a = 2.8 = 2 4/5 [This confirms that your value of a is indeed correct!] y-coordinate = 3a-7 [because N lies on that line] = 7/5 = 1 2/5 so, we proved N = (2 4/5 , 1 2/5) now you have both points Q and N, do you know how to find equation of line given 2 points on it ??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

the equation is

hartnn (hartnn):

also, did you understand my explanation ? ask if any doubts....

OpenStudy (anonymous):

3y+x-7

OpenStudy (anonymous):

No I understood these questions are difficult to me I'm weak in geometry.

OpenStudy (anonymous):

Thanks!

hartnn (hartnn):

did you mean 3y+x=7 ?

hartnn (hartnn):

because thats what i got..

OpenStudy (anonymous):

Well isn't the same ? all I did was carry everything to one side xD

hartnn (hartnn):

ohh..so you didn't mention the =0 part... its correct 3y+x-7=0

OpenStudy (anonymous):

oh sorry :S

hartnn (hartnn):

no, problem and welcome ^_^